OpenStudy (anonymous):

The code to unlock a safety deposit box is three consecutive odd integers whose sum is 117. Find the integers

OpenStudy (anonymous):

I have 29,31,33 just unsure

OpenStudy (anonymous):

So, let the first number be n n + (n+2) + (n+4) = 63 3n + 6 = 63 3n = 63-6 = 57 n = 57/3 = 19 So, the integers are 19,21,23. Let's see if that works. 19+21+23 = 63. Check.